OpenStudy (anonymous):

Given the linear equation 5x−3y=7. If x = 3/5, solve for y. a) y=−3/4 b) y=−4/3 c) y=−10/3 d) y=−3/10 I'm like 90% sure its c

OpenStudy (anonymous):

It is not C

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

The first step is to substitute in 3/5 for x \[5(\frac{ 3 }{ 5 })-3y=7\] Simplify that please

OpenStudy (anonymous):

I cant I don't understand

OpenStudy (anonymous):

Well what is 5 times 3/5?

OpenStudy (mathmale):

You are given an equation in two unknowns, a linear equation. You are told that x=3/5. You need to substitute this value for x in the given equation. This will leave you with an equation in y only (x will have disappeared). Simplify this equation by isolating -3y on the left side of your equation. Finally, solve for y by dividing both sides of the equation by -3. Show whatever work you can do. I'm sure you can handle at least part of this procedure.

OpenStudy (mathmale):

If you don't understand something, please ask for clarification. Being able to do that is a very valuable skill.